Luke is running a fundraising raffle for the local football club. He decides the first prize is $50 more than the second prize, and the third prize half of the second prize. If there is a total of $200 to be used for the prize money, find:
How much each prize winner will recieve.

Answer by nshah11(47)

s = second prize
s + 50 + s + s/2 = 200
s = 150* (2/5) = $60
first prize = $110
third prize = $30